OpenStudy (anonymous):

Thermodynamics: How to solve c) and d) in following scenario: Consider as a system with ideal gas in the cylinder ,the cylinder is fitted with a piston on which a number of small weights are placed. The initial pressure is 200 kPa, and the initial volume of the gas is 0.04 m^3. Will write the problems after posting this.

OpenStudy (anonymous):

a. Let a Bunsen burner be placed under the cylinder, and let the volume of' the gas in- crease to 0.1 m3 while the pressure remains constant. Calculate the work done by the system during this process: Here i got: 200kPa*(0,1m^3-0,04m^3)=12kJ b. Consider the same system and initial conditions, but at the same time as the Bunsen burner is under the cylinder and the piston is rising, let weights be removed from the piston at such a rate that, during the process, the temperature of the gas remains constant. Here i got: (200kPa*0,04m^3)/(0,1m^3)=P2=80kPa c. Consider the same system, but during the heat transfer let the weights be removed at such a rate that the expression PV1. 3 = constant describes the re1ationship between pressure and volume during the process. Again the final volume is 0.1 m3. Calculate the work. d. Consider the system and initial state given in the first three examples, but let the piston be held by a pin so that the volume remains constant. In addition, let heat be transferred from the system until the pressure drops to 100 kPa. Calculate the work. How do i solve c and d?

OpenStudy (anonymous):

And in c: it's PV^1. 3 = constant

OpenStudy (anonymous):

In d) is the answer -4kJ?

OpenStudy (anonymous):

Or is it just 0, since work depends on change in volume?

OpenStudy (anonymous):

what i got so far, which is basicly just guessing but on c: i got (200kPa*(0,04m^3)^1,3)/(0,1m^3)^1,3 = 60,773kPa And d as 100kPa*(0,04m^3-0,04m^3)=0kJ